OpenStudy (goformit100):

Identify the slope and y-intercept of the inequality 9x + 3y > 6. A** slope = 3, y-intercept = 2 B** slope = -3, y-intercept = 2 C** slope = -3, y-intercept = 6 D** slope = negative 1 over 3, y-intercept = 2

OpenStudy (goformit100):

@kelliegirl33

HanAkoSolo (jamierox4ev3r):

9x+3y>6 -9x -9x 3y=-9x+6 /3 /3 /3 y=-3x +2 your answer is B

OpenStudy (anonymous):

we need to put it in y = mx + b form where m is the slope and b is the y intercept. 9x + 3y > 6 3y > -9x + 6 y > -9/3x + 6/3 y > -3x + 2 slope = -3 y intercept = 2

OpenStudy (goformit100):

Thankyou Madam